# Beatty solutions of almost Golomb equations

### 16 Apr 2026 v2

@@ 6.4 A two-letter coarsening
Grouping the three gap values into two classes makes the structure more transparent. Encode $3,4$ as the letter $S$ (short, local gap) and $7$ as the letter $L$ (long gap). The resulting two-letter sequence has frequencies $1-(\sqrt{2}{-}1)=2-\sqrt{2}$ for $S$ and $\sqrt{2}-1$ for $L$. Both frequencies lie in $\mathbb{Q}(\sqrt{2})$, consistent with the fact that $a_{\mathrm{B}}$ is a Sturmian rotation of slope $c=1/\!\sqrt{2}$. Empirically the $S/L$ sequence appears to be Sturmian with slope $\sqrt{2}-1$, but a rigorous identification is left as an open problem.
7 Further remarks
The next remark records the existence of a large family of other nondecreasing solutions for $r=2$.
Remark 22.
The implicit nature of (1) allows multiple self-consistent continuations. For $r=2$, the resulting tree of nondecreasing solutions is a complete binary tree, and a continuum of infinite paths coexist with the greedy and the Beatty. This tree structure, together with an algebraic filter that isolates the $2$-regular paths, is developed in a separate paper [3].
@@ 8 Open problems
    Among nondecreasing solutions of (1) for $r\geq 3$, classify all monotone solutions. For $r=2$ the tree structure is described in [3]; the general case is open.
@@
    Prove Conjecture 21 and identify the substitution (or the primitive system) governing the defect set $\mathcal{D}$ at the right endpoint of the Beatty interval, for $r=2$ and for each $r\geq 3$.

### 21 Apr 2026 v3

@@ 6.3 A two-letter coarsening
Grouping the three gap values into two classes simplifies the structure. Encoding $3,4$ as a short letter $S$ and $7$ as a long letter $L$ yields a two-letter sequence with frequencies $2-\sqrt{2}$ for $S$ and $\sqrt{2}-1$ for $L$. Both frequencies lie in $\mathbb{Q}(\sqrt{2})$, consistent with $a_{\mathrm{B}}$ being a rotation sequence of slope $c=1/\!\sqrt{2}$. Empirically the $S/L$ sequence appears to be Sturmian of slope $\sqrt{2}-1$, but a rigorous identification is left as an open problem.
@@ 7 Open problems
    Among nondecreasing solutions of (1) for $r\geq 3$, classify all monotone solutions. For $r=2$ the tree structure is described in [3]. The general case is open.
@@
    Identify the analogue of Conjecture 26 for each $r\geq 3$. For $r=3$ the defect set at the right endpoint $d=(5\sqrt{3}{-}3)/6$ begins
    $16,\,23,\,42,\,49,\,61,\,68,\,87,\,94,\ldots$
    (A395252), with gap values $\{7,\,12,\,19\}$ corresponding to the consecutive convergent denominators $7$ and $19$ of $1/\!\sqrt{3}$ and their difference. A primitive substitution on $\{7,12,19\}$ should govern the structure, with Perron eigenvalue related to $\sqrt{3}$.